# A novel multipoint stress control volume method for linear elasticity on quadrilateral grids

### v1

@@ 4.2 Quasi-optimal approximation
$(\bm{u},\mathrm{div}\underline{\Pi}_{h}\underline{w})=-(\nabla\bm{u},\underline{\Pi}_{h}\underline{w})=-(\mathcal{A}\underline{\sigma},\underline{\Pi}_{h}\underline{w})-(\text{as}(\underline{\Pi}_{h}\underline{w}),\gamma),$
@@
$(\mathcal{A}\underline{\sigma},\underline{w})+(\bm{u},\mathrm{div}\underline{\Pi}_{h}\underline{w})+(\text{as}(\underline{\Pi}_{h}\underline{w}),\gamma)=(\underline{\sigma},\underline{w}-\underline{\Pi}_{h}\underline{w}),$
which combined with (4.9) implies
$(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h}),\underline{w})+(\bm{u}-\bm{u}_{h},\mathrm{div}\underline{\Pi}_{h}\underline{w})+(\gamma-\gamma_{h},\text{as}(\underline{w}))$ $=(\underline{\sigma},\underline{w}-\underline{\Pi}_{h}\underline{w})\quad\forall\underline{w}\in\underline{\Sigma}_{h}.$ (4.39)
@@
$\begin{split}\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}&\leq C\sup_{\underline{0}\neq\underline{w}\in\underline{\Sigma}_{h}}\frac{(\bm{Q}_{h}\bm{u}-\bm{u}_{h},\mathrm{div}\underline{\Pi}_{h}\underline{w})}{\|\underline{w}\|_{\Sigma,h}}\\ &=C\sup_{\underline{0}\neq\underline{w}\in\underline{\Sigma}_{h}}\frac{-(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h}),\underline{w})+(\underline{\sigma},\underline{w}-\underline{\Pi}_{h}\underline{w})-(\gamma-\gamma_{h},\text{as}(\underline{w}))}{\|\underline{w}\|_{\Sigma,h}}.\end{split}$ (4.40)
@@
$-(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h}),\underline{w})-(\gamma-\gamma_{h},\text{as}(\underline{w}))\leq C\Big{(}\|\underline{\sigma}-\underline{\sigma}_{h}\|_{\mathcal{A}}+\|\gamma-\gamma_{h}\|_{L^{2}(\Omega)}\Big{)}\|\underline{w}\|_{\Sigma,h}.$
@@
$(\underline{\sigma},\underline{w}-\underline{\Pi}_{h}\underline{w})\leq Ch\|\nabla\underline{\sigma}\|_{L^{2}(\Omega)}\|\underline{w}\|_{L^{2}(\Omega)}.$
This and (4.40) imply
$\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}$ $\leq C\Big{(}\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}+\|\gamma-\gamma_{h}\|_{L^{2}(\Omega)}+h|\underline{\sigma}|_{H^{1}(\Omega)}\Big{)},$
which coupled with Corollary 4.1 completes the proof.
@@ 4.3 Superconvergence
$-\mathrm{div}\underline{\psi}$ $=\bm{Q}_{h}\bm{u}-\bm{u}_{h}$ $\quad\mbox{in}\;\Omega,$ (4.41)
$\mathcal{A}\underline{\psi}$ $=\varepsilon(\bm{\phi})$ $\quad\mbox{in}\;\Omega,$ (4.42)
@@
In addition, we let $\underline{\psi}=\nabla\bm{\phi}-\zeta\underline{\delta}$, where $\zeta=\text{rot}(\bm{\phi})$. We let $\bm{\phi}_{h}$, $\underline{\psi}_{h}$ and $\zeta_{h}$ denote the corresponding approximations of $\bm{\phi}$, $\underline{\psi}$ and $\zeta$ based on the discrete formulation (3.2)-(3.4) with the right-hand side $\bm{f}$ replaced by $\bm{Q}_{h}\bm{u}-\bm{u}_{h}$. Then it holds in view of Corollary 4.1
$\|\underline{\psi}-\underline{\psi}_{h}\|_{L^{2}(\Omega)}+\|\zeta-\zeta_{h}\|_{L^{2}(\Omega)}\leq C\Big{(}h|\underline{\psi}|_{H^{1}(\Omega)}+h|\zeta|_{H^{1}(\Omega)}+h\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}\Big{)}.$
@@
(superconvergence). Let $(\underline{\sigma},\bm{u})$ be the exact solution and let $(\underline{\sigma}_{h},\bm{u}_{h},\gamma_{h})\in\underline{\Sigma}_{h}\times\bm{U}_{h}\times\Gamma_{h}$ be the discrete solution of (3.2)-(3.4) obtained by either Method 1 or Method 2. Assume that $\underline{\sigma}\in\underline{H}^{1}(\Omega)$ with $\mathrm{div}\underline{\sigma}\in\bm{H}^{1}(\Omega)$, then the following convergence error estimates hold
$\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}\leq C\Big{(}h^{2}\Big{(}\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\mathrm{div}\underline{\sigma}\|_{H^{1}(\Omega)}\Big{)}+h\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}\Big{)}.$
@@
We multiply (4.41) by $\underline{\sigma}-\underline{\sigma}_{h}$ and (4.42) by $\bm{Q}_{h}\bm{u}-\bm{u}_{h}$, and using the fact that $\underline{\psi}=\nabla\phi-\zeta\underline{\delta}$, we have
$\begin{split}\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}^{2}&=-(\bm{Q}_{h}\bm{u}-\bm{u}_{h},\mathrm{div}\underline{\psi})+(\underline{\psi},\underline{\sigma}-\underline{\sigma}_{h})-(\nabla\bm{\phi},\underline{\sigma}-\underline{\sigma}_{h})\\ &+(\zeta,\text{as}(\underline{\sigma}-\underline{\sigma}_{h}))+(\text{as}(\underline{\psi}),\gamma-\gamma_{h}).\end{split}$ (4.44)
@@
It follows from (4.39) and (4.45) by setting $\underline{w}=\underline{\psi}_{h}$ and $\bm{v}=\bm{\phi}_{h}$ that
$(\underline{\sigma}-\underline{\sigma}_{h},\underline{\psi}_{h})+(\bm{u}-\bm{u}_{h},\mathrm{div}\underline{\Pi}_{h}\underline{\psi}_{h})+(\gamma-\gamma_{h},\text{as}(\underline{\psi}_{h}))$ $=(\underline{\sigma},\underline{\psi}_{h}-\underline{\Pi}_{h}\underline{\psi}_{h})\quad\forall\underline{\psi}_{h}\in\underline{\Sigma}_{h},$ (4.47)
$(\mathrm{div}\underline{\Pi}_{h}(\underline{\sigma}-\underline{\sigma}_{h}),\bm{\phi}_{h})$ $=0\quad\forall\bm{\phi}_{h}\in\bm{U}_{h}.$ (4.48)
As a consequence of (4.47) and (4.44), we have
$\begin{split}\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}^{2}&=-(\bm{Q}_{h}\bm{u}-\bm{u}_{h},\mathrm{div}\underline{\psi})+(\underline{\psi},\underline{\sigma}-\underline{\sigma}_{h})-(\nabla\bm{\phi},\underline{\sigma}-\underline{\sigma}_{h})+(\zeta-\zeta_{h},\text{as}(\underline{\sigma}-\underline{\sigma}_{h}))\\ &=-(\bm{Q}_{h}\bm{u}-\bm{u}_{h},\mathrm{div}(\underline{\psi}-\underline{\Pi}_{h}\underline{\psi}_{h}))+(\underline{\sigma}-\underline{\sigma}_{h},\underline{\psi}-\underline{\psi}_{h})-(\nabla\bm{\phi},\underline{\sigma}-\underline{\sigma}_{h})\\ &+(\underline{\sigma},\underline{\psi}_{h}-\underline{\Pi}_{h}\underline{\psi}_{h})+(\zeta-\zeta_{h},\text{as}(\underline{\sigma}-\underline{\sigma}_{h}))+(\gamma-\gamma_{h},\text{as}(\underline{\psi}-\underline{\psi}_{h})).\end{split}$ (4.49)
@@
$(\underline{\sigma}-\underline{\sigma}_{h},\underline{\psi}-\underline{\psi}_{h})$ $\leq\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}\|\underline{\psi}-\underline{\psi}_{h}\|_{L^{2}(\Omega)}\leq Ch\Big{(}h\|\underline{\sigma}\|_{H^{1}(\Omega)}$
$+\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}\Big{)}\Big{(}\|\underline{\psi}\|_{H^{1}(\Omega)}+\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}\Big{)}.$
The last two terms of (4.49) can be estimated in a similar way. Now we estimate the third term on the right-hand side of (4.49). To this end, an application of integration by parts implies
@@
$=-(\bm{\phi},\mathrm{div}(\underline{\sigma}-\underline{\Pi}_{h}\underline{\sigma}_{h}))+(\underline{\Pi}_{h}\underline{\sigma}_{h}-\underline{\sigma}_{h},\underline{\psi}),$
@@
$(\nabla\bm{\phi},\underline{\sigma}-\underline{\sigma}_{h})$ $=-(\bm{\phi}-\bm{\phi}_{h},\mathrm{div}(\underline{\sigma}-\underline{\Pi}_{h}\underline{\sigma}_{h}))+(\underline{\Pi}_{h}\underline{\sigma}_{h}-\underline{\sigma}_{h},\underline{\psi})$
$=-(\bm{\phi}-\bm{\phi}_{h},\mathrm{div}(\underline{\sigma}-\underline{\Pi}_{h}\underline{\sigma}))+(\underline{\Pi}_{h}\underline{\sigma}_{h}-\underline{\sigma}_{h},\underline{\psi}).$
@@
$(\underline{\Pi}_{h}\underline{\sigma}_{h}-\underline{\sigma}_{h},\underline{\psi})$ $\leq Ch\|\nabla\underline{\psi}\|_{L^{2}(\Omega)}\|\underline{\Pi}_{h}\underline{\sigma}_{h}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}.$
@@
This, (4.50), the interpolation error estimate and Corollary 4.1 yield
$\|\underline{\Pi}_{h}\underline{\sigma}_{h}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}\leq C(h|\underline{\sigma}|_{H^{1}(\Omega)}+h|\gamma|_{H^{1}(\Omega)}+\Big{(}\sum_{M\in\mathcal{T}_{M}}h_{M}^{2}\|\bm{f}\|_{L^{2}(M)}^{2}\Big{)}^{\frac{1}{2}}).$
Similarly, the last term on the right-hand side of (4.49) can be estimated by Lemma 4.8
$(\underline{\sigma},\underline{\psi}_{h}-\underline{\Pi}_{h}\underline{\psi}_{h})$ $\leq Ch\|\nabla\underline{\sigma}\|_{L^{2}(\Omega)}\|\underline{\psi}_{h}-\underline{\Pi}_{h}\underline{\psi}_{h}\|_{L^{2}(\Omega)}$
$\leq Ch^{2}\|\underline{\sigma}\|_{H^{1}(\Omega)}(\|\underline{\psi}\|_{H^{1}(\Omega)}+\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}).$
The proof is completed by using the above estimates and the elliptic regularity estimate (4.43).

### v2

@@ 4.2 The convergence error analysis
$(\bm{u},\mathrm{div}\underline{\Pi}_{h}\underline{w})=-(\nabla\bm{u},\underline{\Pi}_{h}\underline{w})=-(\mathcal{A}\underline{\sigma}+\gamma\underline{\delta},\underline{\Pi}_{h}\underline{w}),$
@@
$(\mathcal{A}\underline{\sigma},\underline{w})+(\bm{u},\mathrm{div}\underline{\Pi}_{h}\underline{w})+(\gamma,\text{as}(w))=(\mathcal{A}\underline{\sigma}+\gamma\underline{\delta},\underline{w}-\underline{\Pi}_{h}\underline{w}),$
which combined with (4.6) implies (4.23).
Then (4.24) follows directly from the definition of $\underline{\Pi}_{h}$ and (4.7). ∎
Theorem 4.2.
Let $(\underline{\sigma},\bm{u})$ be the exact solution and let $(\underline{\sigma}_{h},\bm{u}_{h})\in\underline{\Sigma}_{h}\times\bm{U}_{h}$ be the discrete solution of (4.6)-(4.7). Assume that $(\underline{\sigma},\gamma)\in\underline{H}^{1}(\Omega)\times H^{1}(\Omega)$, then the following convergence error estimates hold
$\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}\leq Ch\Big{(}\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)}\Big{)}$
and
$\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}+\|\gamma_{h}-P_{h}\gamma\|_{L^{2}(\Omega)}\leq Ch\Big{(}\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)}\Big{)}.$
Proof.
Taking $\underline{w}=\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h}$, $\bm{v}=\bm{Q}_{h}\bm{u}-\bm{u}_{h}$ and $\xi=P_{h}\gamma-\gamma_{h}$ in (4.23)-(4.25), then it holds
$(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h}),\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h})+(\bm{u}-\bm{u}_{h},\mathrm{div}\underline{\Pi}_{h}(\underline{I}_{h}^{e}\underline{\sigma}_{h}-\underline{\sigma}_{h}))$ $+(\gamma-\gamma_{h},\text{as}(\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h}))$
$=(\mathcal{A}\underline{\sigma}+\gamma\underline{\delta},\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h}-\underline{\Pi}_{h}(\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h})),$
$(\mathrm{div}\underline{\Pi}_{h}(\underline{\sigma}-\underline{\sigma}_{h}),\bm{Q}_{h}\bm{u}-\bm{u}_{h})$ $=0,$
$(\text{as}(\underline{\sigma}-\underline{\sigma}_{h}),P_{h}\gamma-\gamma_{h})$ $=0.$
The definition of $\underline{\Pi}_{h}$ (cf. (4.3)) yields
$(\mathrm{div}\underline{\Pi}_{h}(\underline{\sigma}-\underline{\sigma}_{h}),\bm{Q}_{h}\bm{u}-\bm{u}_{h})=(\mathrm{div}(\underline{\sigma}-\underline{\Pi}_{h}\underline{\sigma}_{h}),\bm{Q}_{h}\bm{u}-\bm{u}_{h}).$
We can infer from the definition of $\underline{\Pi}_{h}$ (cf. (4.3)), integration by parts and (4.2) that
$(\bm{u}-\bm{u}_{h},\mathrm{div}\underline{\Pi}_{h}(\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h}))$ $=(\bm{Q}_{h}\bm{u}-\bm{u}_{h},\mathrm{div}\underline{\Pi}_{h}(\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h}))$
$=\sum_{T\in\mathcal{T}_{M}}(\bm{Q}_{h}\bm{u}-\bm{u}_{h},\underline{\Pi}_{h}(\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h})\bm{n})_{\partial T}$
$=\sum_{T\in\mathcal{T}_{M}}(\bm{Q}_{h}\bm{u}-\bm{u}_{h},(\underline{I}_{h}^{e}\underline{\sigma}-\underline{\Pi}_{h}\underline{u}_{h})\bm{n})_{\partial T}$
$=\sum_{T\in\mathcal{T}_{M}}(\bm{Q}_{h}\bm{u}-\bm{u}_{h},(\underline{\sigma}-\underline{\Pi}_{h}\underline{\sigma}_{h})\bm{n})_{\partial T}$
$=(\bm{Q}_{h}\bm{u}-\bm{u}_{h},\mathrm{div}(\underline{\sigma}-\underline{\Pi}_{h}\underline{\sigma}_{h})).$
In view of Lemma 4.4 and the Piola transformation, we have
$(\mathcal{A}\underline{\sigma}+\gamma\underline{\delta},\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h}-\underline{\Pi}_{h}(\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h}))$ $\leq Ch(\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)})\|\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}.$
Therefore, we have from the Cauchy-Schwarz inequality that
$\begin{split}\|\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}^{2}&=(\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma},\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h})+(\mathcal{A}\underline{\sigma}+\gamma\underline{\delta},\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h}-\underline{\Pi}_{h}(\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h}))\\ &-(\gamma-P_{h}\gamma,\text{as}(\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h}))-(\text{as}(\underline{\sigma}-\underline{I}_{h}^{e}\underline{\sigma}),P_{h}\gamma-\gamma_{h})\\ &\leq C\Big{(}h(\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)})+\|P_{h}\gamma-\gamma_{h}\|_{L^{2}(\Omega)}\Big{)}\|\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}.\end{split}$ (4.26)
@@
$\begin{split}\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}+\|\gamma_{h}-P_{h}\gamma\|_{L^{2}(\Omega)}&\leq C\sup_{\underline{0}\neq\underline{w}\in\underline{\Sigma}_{h}}\frac{(\bm{Q}_{h}\bm{u}-\bm{u}_{h},\mathrm{div}\underline{\Pi}_{h}\underline{w})+(\text{as}(\underline{w}),P_{h}\gamma-\gamma_{h})}{\|\underline{w}\|_{L^{2}(\Omega)}}\\ &=C\sup_{\underline{0}\neq\underline{w}\in\underline{\Sigma}_{h}}\frac{-(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h}),\underline{w})+(\mathcal{A}\underline{\sigma}+\gamma\underline{\delta},\underline{w}-\underline{\Pi}_{h}\underline{w})-(\gamma-P_{h}\gamma,\text{as}(\underline{w}))}{\|\underline{w}\|_{L^{2}(\Omega)}}.\end{split}$ (4.27)
@@
$-(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h}),\underline{w})-(\gamma-P_{h}\gamma,\text{as}(\underline{w}))\leq C\Big{(}\|\underline{\sigma}-\underline{\sigma}_{h}\|_{\mathcal{A}}+\|\gamma-P_{h}\gamma\|_{L^{2}(\Omega)}\Big{)}\|\underline{w}\|_{L^{2}(\Omega)}.$
@@
$(\mathcal{A}\underline{\sigma}+\gamma\underline{\delta},\underline{w}-\underline{\Pi}_{h}\underline{w})\leq Ch(\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)})\|\underline{w}\|_{L^{2}(\Omega)}.$
Therefore, we have
$\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}+\|\gamma_{h}-P_{h}\gamma\|_{L^{2}(\Omega)}$ $\leq C\Big{(}\|\underline{\sigma}-\underline{\sigma}_{h}\|_{\mathcal{A}}+\|\gamma-P_{h}\gamma\|_{L^{2}(\Omega)}+h(\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)})\Big{)}.$ (4.28)
Combining the above with (4.26) implies
$\|\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}\leq Ch(\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)}).$
As a consequence of the triangle inequality and the interpolation error estimate, one has
$\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}\leq\|\underline{\sigma}-\underline{I}_{h}^{e}\underline{\sigma}\|_{L^{2}(\Omega)}+\|\underline{I}_{h}^{e}\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}\leq Ch(\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)}).$
This and (4.28) give
$\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}+\|\gamma_{h}-P_{h}\gamma\|_{L^{2}(\Omega)}$ $\leq Ch(\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)}).$
Therefore, the proof is completed.
@@ 4.3 Superconvergence
$\mathrm{div}\underline{\psi}$ $=\bm{Q}_{h}\bm{u}-\bm{u}_{h}$ $\quad\mbox{in}\;\Omega,$ (4.29)
$\mathcal{A}\underline{\psi}$ $=-\varepsilon(\bm{\phi})$ $\quad\mbox{in}\;\Omega,$ (4.30)
@@
In addition, we let $\mathcal{A}\underline{\psi}=-\nabla\bm{\phi}+\zeta\underline{\delta}$, where $\zeta=\text{rot}(\bm{\phi})$. We let $\bm{\phi}_{h}$, $\underline{\psi}_{h}$ and $\zeta_{h}$ denote the corresponding approximations of $\bm{\phi}$, $\underline{\psi}$ and $\zeta$ based on the discrete formulation (3.2)-(3.4) with the right-hand side $\bm{f}$ replaced by $\bm{Q}_{h}\bm{u}-\bm{u}_{h}$. Then it holds in view of Theorem 4.2 and (4.31)
$\begin{split}\|\underline{\psi}-\underline{\psi}_{h}\|_{L^{2}(\Omega)}+\|\zeta-\zeta_{h}\|_{L^{2}(\Omega)}&\leq C\Big{(}h\|\underline{\psi}\|_{H^{1}(\Omega)}+h\|\zeta\|_{H^{1}(\Omega)}+h\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}\Big{)}\\ &\leq Ch\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}.\end{split}$ (4.32)
@@
(superconvergence). Let $(\underline{\sigma},\bm{u})$ be the exact solution and let $(\underline{\sigma}_{h},\bm{u}_{h},\gamma_{h})\in\underline{\Sigma}_{h}\times\bm{U}_{h}\times\Gamma_{h}$ be the discrete solution of (3.2)-(3.4) obtained by either Method 1 or Method 2. Assume that $\underline{\sigma}\in\underline{H}^{1}(\Omega)$ with $\mathrm{div}\underline{\sigma}\in\bm{H}^{1}(\Omega)$ and $\gamma\in H^{1}(\Omega)$, then the following convergence error estimate holds
$\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}\leq Ch^{2}\Big{(}\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)}+\|\mathrm{div}\underline{\sigma}\|_{H^{1}(\Omega)}\Big{)}.$
@@
We multiply (4.29) by $\underline{\sigma}-\underline{\sigma}_{h}$ and (4.30) by $\bm{Q}_{h}\bm{u}-\bm{u}_{h}$, and using the fact that $\mathcal{A}\underline{\psi}=-\nabla\bm{\phi}+\zeta\underline{\delta}$, we have
$\begin{split}\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}^{2}&=(\bm{Q}_{h}\bm{u}-\bm{u}_{h},\mathrm{div}\underline{\psi})+(\mathcal{A}\underline{\psi},\underline{\sigma}-\underline{\sigma}_{h})+(\nabla\bm{\phi},\underline{\sigma}-\underline{\sigma}_{h})\\ &-(\zeta,\text{as}(\underline{\sigma}-\underline{\sigma}_{h}))+(\text{as}(\underline{\psi}),\gamma-\gamma_{h}).\end{split}$ (4.33)
@@
It follows from (4.23)-(4.25) by setting $\underline{w}=\underline{\psi}_{h}$, $\bm{v}=\bm{\phi}_{h}$ and $\xi=\zeta_{h}$ that
$(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h}),\underline{\psi}_{h})+(\bm{u}-\bm{u}_{h},\mathrm{div}\underline{\Pi}_{h}\underline{\psi}_{h})+(\gamma-\gamma_{h},\text{as}(\underline{\psi}_{h}))$ $=(\mathcal{A}\underline{\sigma}+\gamma\underline{\delta},\underline{\psi}_{h}-\underline{\Pi}_{h}\underline{\psi}_{h}),$ (4.36)
$(\mathrm{div}\underline{\Pi}_{h}(\underline{\sigma}-\underline{\sigma}_{h}),\bm{\phi}_{h})$ $=0.$ (4.37)
$(\textnormal{as}(\underline{\sigma}-\underline{\sigma}_{h}),\zeta_{h})$ $=0.$ (4.38)
As a consequence of (4.36)-(4.38) and (4.33), we have
$\begin{split}\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}^{2}&=(\bm{Q}_{h}\bm{u}-\bm{u}_{h},\mathrm{div}\underline{\psi})+(\mathcal{A}\underline{\psi},\underline{\sigma}-\underline{\sigma}_{h})+(\nabla\bm{\phi},\underline{\sigma}-\underline{\sigma}_{h})-(\zeta-\zeta_{h},\text{as}(\underline{\sigma}-\underline{\sigma}_{h}))\\ &+(\text{as}(\underline{\psi}),\gamma-\gamma_{h})\\ &=(\bm{Q}_{h}\bm{u}-\bm{u}_{h},\mathrm{div}(\underline{\psi}-\underline{\Pi}_{h}\underline{\psi}_{h}))+(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h}),\underline{\psi}-\underline{\psi}_{h})+(\nabla\bm{\phi},\underline{\sigma}-\underline{\sigma}_{h})\\ &+(\mathcal{A}\underline{\sigma}+\gamma\underline{\delta},\underline{\psi}_{h}-\underline{\Pi}_{h}\underline{\psi}_{h})-(\zeta-\zeta_{h},\text{as}(\underline{\sigma}-\underline{\sigma}_{h}))+(\gamma-\gamma_{h},\text{as}(\underline{\psi}-\underline{\psi}_{h})).\end{split}$ (4.39)
@@
$(\mathcal{A}(\underline{\sigma}-\underline{\sigma}_{h}),\underline{\psi}-\underline{\psi}_{h})$ $\leq\|\underline{\sigma}-\underline{\sigma}_{h}\|_{\mathcal{A}}\|\underline{\psi}-\underline{\psi}_{h}\|_{L^{2}(\Omega)}\leq Ch^{2}\Big{(}\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)}\Big{)}\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}.$
Now we estimate the third term on the right-hand side of (4.39). To this end, an application of integration by parts implies
@@
$=-(\bm{\phi},\mathrm{div}(\underline{\sigma}-\underline{\Pi}_{h}\underline{\sigma}_{h}))-(\underline{\Pi}_{h}\underline{\sigma}_{h}-\underline{\sigma}_{h},\mathcal{A}\underline{\psi}-\zeta\underline{\delta}),$
@@
$(\nabla\bm{\phi},\underline{\sigma}-\underline{\sigma}_{h})$ $=-(\bm{\phi}-\bm{\phi}_{h},\mathrm{div}(\underline{\sigma}-\underline{\Pi}_{h}\underline{\sigma}_{h}))-(\underline{\Pi}_{h}\underline{\sigma}_{h}-\underline{\sigma}_{h},\mathcal{A}\underline{\psi}-\zeta\underline{\delta})$
$=-(\bm{\phi}-\bm{\phi}_{h},\mathrm{div}(\underline{\sigma}-\underline{\Pi}_{h}\underline{\sigma}))-(\underline{\Pi}_{h}\underline{\sigma}_{h}-\underline{\sigma}_{h},\mathcal{A}\underline{\psi}-\zeta\underline{\delta}).$
@@
$(\underline{\Pi}_{h}\underline{\sigma}_{h}-\underline{\sigma}_{h},\mathcal{A}\underline{\psi}-\zeta\underline{\delta})$ $\leq Ch(\|\underline{\psi}\|_{H^{1}(\Omega)}+\|\zeta\|_{H^{1}(\Omega)})\|\underline{\Pi}_{h}\underline{\sigma}_{h}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}.$
@@
This, (4.40), the interpolation error estimate and Theorem 4.2 yield
$\|\underline{\Pi}_{h}\underline{\sigma}_{h}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}\leq C(h\|\underline{\sigma}\|_{H^{1}(\Omega)}+h\|\gamma\|_{H^{1}(\Omega)}).$
Similarly, the fourth term on the right-hand side of (4.39) can be estimated by Lemma 4.5
$(\mathcal{A}\underline{\sigma}+\gamma\underline{\delta},\underline{\psi}_{h}-\underline{\Pi}_{h}\underline{\psi}_{h})$ $\leq Ch(\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)})\|\underline{\psi}_{h}-\underline{\Pi}_{h}\underline{\psi}_{h}\|_{L^{2}(\Omega)}$
$\leq Ch^{2}(\|\underline{\sigma}\|_{H^{1}(\Omega)}+\|\gamma\|_{H^{1}(\Omega)})\|\bm{Q}_{h}\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}.$
The last two terms on the right-hand side of (4.39) can be bounded by the Cauchy-Schwarz inequality, Theorem 4.2 and (4.32). The proof is completed by using the above estimates.